# MMLongBench-Doc: Benchmarking Long-context Document Understanding with Visualizations

### v2

@@ Abstract
Understanding documents with rich layouts and multi-modal components is a long-standing and practical task. Recent Large Vision-Language Models (LVLMs) have made remarkable strides in various tasks, particularly in single-page document understanding (DU). However, their abilities on long-context DU remain an open problem. This work presents MMLongBench-Docfootnotetext: Project Page: https://mayubo2333.github.io/MMLongBench-Doc, a long-context, multi-modal benchmark comprising 1,091 expert-annotated questions. Distinct from previous datasets, it is constructed upon 135 lengthy PDF-formatted documents with an average of 47.5 pages and 21,214 textual tokens. Towards comprehensive evaluation, answers to these questions rely on pieces of evidence from (1) different sources (text, image, chart, table, and layout structure) and (2) various locations (i.e., page number). Moreover, 33.0% of the questions are cross-page questions requiring evidence across multiple pages. 22.5% of the questions are designed to be unanswerable for detecting potential hallucinations. Experiments on 14 LVLMs demonstrate that long-context DU greatly challenges current models. Notably, the best-performing model, GPT-4o, achieves an F1 score of only 44.9%, while the second-best, GPT-4V, scores 30.5%. Furthermore, 12 LVLMs (all except GPT-4o and GPT-4V) even present worse performance than their LLM counterparts which are fed with lossy-parsed OCR documents. These results validate the necessity of future research toward more capable long-context LVLMs.
@@ 1 Introduction
In this paper, we present MMLongBench-Doc, a benchmark designed to evaluate the Multi-Modality Long-context Document understanding abilities of LVLMs. Towards a comprehensive benchmark, it incorporates lengthy documents from both four existing datasets [13; 17; 18; 19] and other various papers, brochures, etc. Consequently, our benchmark includes 135 PDF-formatted documents spanning across 7 diverse domains, with each document averaging 47.5 pages and 21,214.1 textual tokens. Regarding the questions, we employ ten expert-level annotators to (1) edit questions associated with documents from existing datasets to meet our benchmark’s standard and (2) create new questions for all collected documents to expand the scale of the benchmark. Then a three-round, semi-automatic reviewing process ensures the benchmark’s annotation quality. As a result, MMLongBench-Doc comprises 1,091 human-annotated questions, with 184 sourced from four existing datasets and 907 newly annotated. Being a multi-modal benchmark, the answer to each question requires evidence from one or more of these five in-document sources: text, layout, chart, table, and image. Questions are categorized into three types based on the number of evidence pages  [^1] Given a document $D$ and a question $q$ upon $D$, We call page $P$ (in document $D$) an evidence page of $q$ if the answer of $q$ necessitates one or more pieces of evidence in page $P$., with examples illustrated in Figure 1(a): (1) 485 single-page questions (with one evidence page) mainly to evaluate localization abilities, (2) 360 cross-page questions (with multiple evidence pages) to assess cross-page comprehension, and (3) 246 unanswerable questions (no evidence for answering it, i.e., no evidence pages) to reduce shortcuts and measure LVLMs’ potential hallucinations. Meta-information including evidence pages, sources, and answer formats, is preserved for fine-grained evaluation and analysis. Detailed descriptions of the annotation pipeline and statistics can be found in Section 3.
@@ 3 MMLongBench-Doc
As illustrated in Figure 2, we design a three-stage annotation pipeline for the construction of our benchmark. The three stages (document collection, question&answer collection, and quality control) will be introduced in Section 3.1, Section 3.2, and Section 3.3, respectively. We also provide key statistics of our benchmark in Section 3.4.
Figure 2: The annotation pipeline of MMLongBench-Doc.
@@ 3.1 Document Collection
In summary, we collect a total of 135 documents. Among them, 76 documents are from existing datasets (colored in purple and exemplified as Document 2 in Figure 2, the same below) and incorporate previously annotated questions (represented as triangles). The remaining 59 documents are newly collected (colored in orange and exemplified as Document 1) and incorporate no existing questions. We manually categorize them into 7 types: Research Report, Financial Report, Academic Paper, Brochure, Guideline, Administration & Industry File, Tutorial / Workshop. We showcase some instances of these documents in Appendix A.3.
@@ 3.2 Question and Answer Collection
We employ ten expert-level annotators, including some authors, for the question-and-answer collection. All of them are doctors or Ph.D. students proficient in English reading and writing. Before formal annotation, they undergo a training session and pre-annotate three documents for practice. We iteratively review their annotation results and provide personalized feedback until their annotations meet the standards mentioned above. Regarding the formal annotation, we divide 135 documents into 54 batches (each having 2-4 documents) and dispatch these batches to annotators. We then ask the annotators to submit their results in units of batches and set reasonable time intervals for each batch’s submission. We timely evaluate their annotations after each submission and remind the annotators if their questions in this turn diverge from the standards. It avoids the annotators rushing all assignments in a short time and benefits the annotation quality. We recommend the annotators take 60-90 minutes on each document. Specifically, the annotators shall rapidly read through the whole document in the first 15-30 minutes. For the remaining time, they shall dive deep into specific components to modify existing annotations and/or add new annotations as detailed below.
@@
Specifically, we classify six potential problems in original annotations: Wrong Answers or Evidence Pages, Repetitive Question, Ambiguous Question, Decontextualization-required Question, Low Document-relevant Question and Potential Shortcut. See detailed explanations and examples about these problems in Appendix A.4. Given an existing document, the annotators are tasked to evaluate each existing question’s quality according to whether they have one or more above problems and assign a label from {Retain, Revise, Remove} for each question. Then the annotators would revise the Revise questions to meet our quality criteria and remove the Remove questions. See triangle-marked questions colored in green, yellow, and red in Figure 2 as Retain, Revise, Remove examples, respectively. Among all 425 original questions from 76 existing documents, 32.2% of them are revised and 46.1% are removed. We finally collect 211 questions in this procedure. The corresponding GUI is shown in Appendix A.6.
Add New Questions. We newly annotate questions on both existing and newly collected documents to expand the questions in our benchmark. See square-marked questions in Figure 2 as examples. Specifically, we ask annotators to add about 3 questions on existing documents, and 6 questions on newly-collected documents. Given most existing questions (even after editing) are single-page ones and sourced from texts, we put more focus on (1) cross-page and unanswerable questions and (2) questions sourced from tables, charts, and images for newly added questions to balance the distribution. We detail the quantitative requirements in Appendix A.5. Associated with questions, annotators also provide reference answers and meta-information (i.e., evidence sources, answer format, evidence locations) for all samples. We finalized a collection of 965 samples in this procedure. The corresponding GUI is shown in Appendix A.6.
@@ 3.3 Quality Control
Combining the merits of both human annotators and cutting-edge LVLMs, we adopt a three-round, semi-automatic quality control procedure to further improve the annotation quality of our benchmark. We detail each round in the following components.
Document-relevant Detection. Our benchmark is designed to evaluate LVLMs’ long-context document understanding abilities. All questions are expected to be unanswerable without access to corresponding documents. To remove low document-relevant questions (i.e., questions not relying on documents), we feed each annotated question WITHOUT documents to GPT-4o. A question will be identified as low document-relevant question if GPT-4o correctly predicts under this case. Ultimately, 85 samples are identified as low document-relevant questions and removed in this round.
@@
Total questions 1,091
- Single-page question 485 (44.5%)
- Cross-page questions 360 (33.0%)
- Unanswerable questions 246 (22.5%)
- Derived questions 184 (16.9%)
- Newly-annotated questions 907 (83.1%)
@@
- Pure-text 298 (35.3%)
- Layout 119 (14.1%)
- Table 218 (25.8%)
- Chart 178 (21.1%)
- Image 293 (34.7%)
@@
- String 250 (29.5%)
- Integer 290 (34.2%)
- Float 160 (18.9%)
- List 147 (17.3%)
@@ 3.4 Dataset Overview and Analysis
The main statistics of MMLongBench-Doc are presented in Table 2. Overall, our benchmark consists of 1,091 questions. These questions are constructed upon 135 lengthy documents across 7 document types, with an average of 47.5 pages and 21,214.1 tokens. Please see detailed distributions of these documents in Figure 3. Regarding the questions, there are 485 single-page questions (1 evidence page), 360 cross-page questions (2+ evidence pages), and 246 unanswerable questions (no evidence page). These three types of questions evaluate the LVLMs’s long-context DU capabilities from complementary aspects: the localization ability, the cross-page comprehension ability, and the hallucination severity, respectively. For single-page and cross-page questions, their answer evidence is scattered among different context sources (i.e., text, layout, table, chart, image) and evenly distributed across different locations of the documents (see Table 2, Figure 4 Left and Middle). Also notably, 29.0% of cross-page questions have more than two evidence pages, which further enhances the challenge of our benchmark.
4 Evaluation
4.1 Evaluation Protocol
4.2 Experimental Setup
For comparison, we also use the Tesseract [43] OCR model to recognize and extract texts from the documents and feed the parsed documents to 10 LLMs, including 6 proprietary and 4 open-source ones. Texts exceeding their context lengths are truncated. Notably, as a key component of the classical solution for the DU task, the OCR model can handle most flattened texts and some structured tables in the document. However, it cannot perceive the information from the charts or images. Thus the TXT-formatted, OCR-parsed documents are lossy documents in which the information is not fully preserved. More detailed hyperparameters are introduced in Appendix B.5.
@@ 4.3 Main Results
We compare the performance of different LVLMs and LLMs in Table 3, reporting their generalized accuracy and F1 scores (shown in the last two columns). Regarding LVLMs, we draw several conclusions as below: (1) The performance demonstrates that long-context DU is still a challenging and unsolved task for current LVLMs. The best-performing LVLM, GPT-4o, merely achieves a 44.9% F1 score. The second best-performing LVLM, GPT-4V, lags behind by over 10% percent and presents a 31.4% F1 score. All other LVLMs only achieve about 20% or even lower F1 scores. (2) Though far from satisfactory, GPT-4o performs much better than all other models (including GPT-4V). Thus we speculate that the multi-modal pre-training paradigm significantly benefits LVLMs’ cross-modality understanding capabilities. (3) Proprietary LVLMs perform better than open-source LVLMs by a large margin. We attribute it to the difference of acceptable image numbers: open-source LVLMs only support single-image or several-image  [^6] We observe that all open-source LVLMs struggle when fed with more than 5 images. inputs, while proprietary LVLMs can be fed with at least 20 images or even more. Given that lengthy documents have tens of even hundreds of pages, it is impractical for open-source LVLMs to accurately perceive the information in the documents from the excessively concatenated images. (4) The performances of different models are highly correlated with their acceptable image numbers and maximum image resolutions. Notably, open-source LVLMs that support high-resolution images (i.e., InternLM-XC2-4KHD and InternVL-Chat-v1.5) exhibit superior performance compared to those with lower resolution limits.
@@ 5.2 Error Analysis
5.3 Case Study
In this section, we compare the responses from six representative models, comprising four LVLMs (GPT-4o, GPT-4V, Gemini-1.5-Pro, InternVL-Chat-v1.5) and two LLMs (GPT-4 and Mixtral 8x22B). We show one case in Figure 9 and four more cases in Appendix C.2. It shows that: (1) GPT-4o outperforms all other models by a significant margin. (2) While LLMs demonstrate performance comparable to LVLMs on text-related questions (see Figure 38), they lag behind on questions concerning other vision-intensive sources such as charts, layouts, and images (see Figures 9, 39, and 40). (3) Gemini-Pro-1.5 tends to adopt a conservative approach, often opting not to provide an answer (see Figures 9, 37, 38, and 39). (4) Even the best-performing open-source LVLM, InternVL-v1.5, struggles with these cases, indicating that open-source LVLMs still fall short in tasks involving long-context document understanding.
Figure 9: Case Study. Evidence sources: two charts and one table. The three evidence pages of this question are zoomed in. We manually analyze the responses step-by-step: the correct extracted information and reasoning are colored in green, and the wrong ones are colored in red.
@@ A.5 New Question Annotation
A.6 GUI Screenshots

### v3

@@ Abstract
Understanding documents with rich layouts and multi-modal components is a long-standing and practical task. Recent Large Vision-Language Models (LVLMs) have made remarkable strides in various tasks, particularly in single-page document understanding (DU). However, their abilities on long-context DU remain an open problem. This work presents MMLongBench-Docfootnotetext: Project Page: https://mayubo2333.github.io/MMLongBench-Doc, a long-context, multi-modal benchmark comprising 1,082 expert-annotated questions. Distinct from previous datasets, it is constructed upon 135 lengthy PDF-formatted documents with an average of 47.5 pages and 21,214 textual tokens. Towards comprehensive evaluation, answers to these questions rely on pieces of evidence from (1) different sources (text, image, chart, table, and layout structure) and (2) various locations (i.e., page number). Moreover, 33.7% of the questions are cross-page questions requiring evidence across multiple pages. 20.6% of the questions are designed to be unanswerable for detecting potential hallucinations. Experiments on 14 LVLMs demonstrate that long-context DU greatly challenges current models. Notably, the best-performing model, GPT-4o, achieves an F1 score of only 44.9%, while the second-best, GPT-4V, scores 30.5%. Furthermore, 12 LVLMs (all except GPT-4o and GPT-4V) even present worse performance than their LLM counterparts which are fed with lossy-parsed OCR documents. These results validate the necessity of future research toward more capable long-context LVLMs.
@@ 1 Introduction
In this paper, we present MMLongBench-Doc, a benchmark designed to evaluate the Multi-Modality Long-context Document understanding abilities of LVLMs. Towards a comprehensive benchmark, it incorporates lengthy documents from both four existing datasets [13; 17; 18; 19] and other various papers, brochures, etc. Consequently, our benchmark includes 135 PDF-formatted documents spanning across 7 diverse domains, with each document averaging 47.5 pages and 21,214.1 textual tokens. Regarding the questions, we employ ten expert-level annotators to (1) edit questions associated with documents from existing datasets to meet our benchmark’s standard and (2) create new questions for all collected documents to expand the scale of the benchmark. Then a three-round, semi-automatic reviewing process ensures the benchmark’s annotation quality. As a result, MMLongBench-Doc comprises 1,082 human-annotated questions, with 184 sourced from four existing datasets and 898 newly annotated. Being a multi-modal benchmark, the answer to each question requires evidence from one or more of these five in-document sources: text, layout, chart, table, and image. Questions are categorized into three types based on the number of evidence pages  [^1] Given a document $D$ and a question $q$ upon $D$, We call page $P$ (in document $D$) an evidence page of $q$ if the answer of $q$ necessitates one or more pieces of evidence in page $P$., with examples illustrated in Figure 1(a): (1) 494 single-page questions (with one evidence page) mainly to evaluate localization abilities, (2) 365 cross-page questions (with multiple evidence pages) to assess cross-page comprehension, and (3) 223 unanswerable questions (no evidence for answering it, i.e., no evidence pages) to reduce shortcuts and measure LVLMs’ potential hallucinations. Meta-information including evidence pages, sources, and answer formats, is preserved for fine-grained evaluation and analysis. Detailed descriptions of the annotation pipeline and statistics can be found in Section 3.
@@ 3 MMLongBench-Doc
We design a three-stage annotation pipeline for the construction of our benchmark. The three stages will be introduced in Section 3.1, Section 3.2, and Section 3.3, respectively. We also provide key statistics of our benchmark in Section 3.4.
@@ 3.1 Document Collection
In summary, we collect a total of 135 documents. Among them, 76 documents are from existing datasets and incorporate previously annotated questions (represented as triangles). The remaining 59 documents are newly collected and incorporate no existing questions. We manually categorize them into 7 types: Research Report, Financial Report, Academic Paper, Brochure, Guideline, Administration & Industry File, Tutorial / Workshop. We showcase some instances of these documents in Appendix A.3.
@@ 3.2 Question and Answer Collection
Ten authors serve as expert-level annotators for the question-and-answer collection. All of them are doctors or Ph.D. students proficient in English reading and writing. Before formal annotation, they undergo a training session and pre-annotate three documents for practice. We iteratively review their annotation results and provide personalized feedback until their annotations meet the standards mentioned above. Regarding the formal annotation, we divide 135 documents into 54 batches (each having 2-4 documents) and dispatch these batches to annotators. We then ask the annotators to submit their results in units of batches and set reasonable time intervals for each batch’s submission. We timely evaluate their annotations after each submission and remind the annotators if their questions in this turn diverge from the standards. It avoids the annotators rushing all assignments in a short time and benefits the annotation quality. We recommend the annotators take 60-90 minutes on each document. Specifically, the annotators shall rapidly read through the whole document in the first 15-30 minutes. For the remaining time, they shall dive deep into specific components to modify existing annotations and/or add new annotations as detailed below.
@@
Specifically, we classify six potential problems in original annotations: Wrong Answers or Evidence Pages, Repetitive Question, Ambiguous Question, Decontextualization-required Question, Low Document-relevant Question and Potential Shortcut. See detailed explanations and examples about these problems in Appendix A.4. Given an existing document, the annotators are tasked to evaluate each existing question’s quality according to whether they have one or more above problems and assign a label from {Retain, Revise, Remove} for each question. Then the annotators would revise the Revise questions to meet our quality criteria and remove the Remove questions. Among all 425 original questions from 76 existing documents, 32.2% of them are revised and 46.1% are removed. We finally collect 211 questions in this procedure. The corresponding GUI is shown in Appendix A.7.
Add New Questions. We newly annotate questions on both existing and newly collected documents to expand the questions in our benchmark. Specifically, we ask annotators to add about 3 questions on existing documents, and 6 questions on newly-collected documents. Given most existing questions (even after editing) are single-page ones and sourced from texts, we put more focus on (1) cross-page and unanswerable questions and (2) questions sourced from tables, charts, and images for newly added questions to balance the distribution. We detail the quantitative requirements in Appendix A.5. Associated with questions, annotators also provide reference answers and meta-information (i.e., evidence sources, answer format, evidence locations) for all samples. We finalized a collection of 965 samples in this procedure. The corresponding GUI is shown in Appendix A.7.
@@ 3.3 Quality Control
Combining the merits of humans and LVLMs, we adopt a three-round, semi-automatic quality control procedure to improve the annotation quality of our benchmark. We detail each round in the following components and leave the discussion of potential bias in Appendix A.6.
Document-relevant Detection. Our benchmark is designed to evaluate LVLMs’ long-context document understanding abilities. All questions are expected to be unanswerable without access to corresponding documents. To remove low document-relevant questions (i.e., questions not relying on documents), we feed each annotated question WITHOUT documents to GPT-4o. A question will be identified as low document-relevant question if GPT-4o correctly predicts under this case. Ultimately, 94 samples are identified as low document-relevant questions and removed in this round.
@@
Total questions 1,082
- Single-page question 494 (45.7%)
- Cross-page questions 365 (33.7%)
- Unanswerable questions 223 (20.6%)
- Derived questions 184 (17.0%)
- Newly-annotated questions 898 (83.0%)
@@
- Pure-text 305 (35.5%)
- Layout 119 (13.9%)
- Table 218 (25.4%)
- Chart 178 (20.7%)
- Image 304 (35.4%)
@@
- String 250 (29.1%)
- Integer 299 (34.8%)
- Float 159 (18.5%)
- List 151 (17.6%)
@@ 3.4 Dataset Overview and Analysis
The main statistics of MMLongBench-Doc are presented in Table 2. Overall, our benchmark consists of 1,082 questions. These questions are constructed upon 135 lengthy documents across 7 document types, with an average of 47.5 pages and 21,214.1 tokens. Please see detailed distributions of these documents in Figure 2. Regarding the questions, there are 494 single-page questions (1 evidence page), 365 cross-page questions (2+ evidence pages), and 223 unanswerable questions (no evidence page). These three types of questions evaluate the LVLMs’s long-context DU capabilities from complementary aspects: the localization ability, the cross-page comprehension ability, and the hallucination severity, respectively. For single-page and cross-page questions, their answer evidence is scattered among different context sources (i.e., text, layout, table, chart, image) and evenly distributed across different locations of the documents (see Table 2, Figure 3 Left and Middle). Also notably, 28.6% of cross-page questions have more than two evidence pages, which further enhances the challenge of our benchmark.
@@
Human Baseline
Human Experts - - - - - - - - - - 65.8 66.0
4 Evaluation
4.1 Evaluation Protocol
4.2 Experimental Setup
For comparison, we also use the Tesseract [42] OCR model to recognize and extract texts from the documents and feed the parsed documents to 10 LLMs, including 6 proprietary and 4 open-source ones. Texts exceeding their context lengths are truncated. Notably, as a key component of the classical solution for the DU task, the OCR model can handle most flattened texts and some structured tables in the document. However, it cannot perceive the information from the charts or images. Thus the TXT-formatted, OCR-parsed documents are lossy documents in which the information is not fully preserved. More detailed hyperparameters are introduced in Appendix B.5. Additionally, we also conduct manual evaluation on a subset of our datasets (238 questions from 29 documents) to indicate the difficulty of this task for humans.
@@ 4.3 Main Results
We compare the performance of different LVLMs and LLMs in Table 3, reporting their generalized accuracy and F1 scores (shown in the last two columns). Regarding LVLMs, we draw several conclusions as below: (1) The performance demonstrates that long-context DU is still a challenging and unsolved task for current LVLMs. The best-performing LVLM, GPT-4o, merely achieves a 44.9% F1 score. The second best-performing LVLM, GPT-4V, lags behind by over 10% percent and presents a 31.4% F1 score. All other LVLMs only achieve about 20% or even lower F1 scores. (2) Though far from satisfactory, GPT-4o performs much better than all other models (including GPT-4V). Thus we speculate that the multi-modal pre-training paradigm significantly benefits LVLMs’ cross-modality understanding capabilities. (3) Proprietary LVLMs perform better than open-source LVLMs by a large margin. We attribute it to the difference of acceptable image numbers: open-source LVLMs only support single-image or several-image inputs, while proprietary LVLMs can be fed with at least 20 images or even more. Given that lengthy documents have tens of even hundreds of pages, it is impractical for open-source LVLMs to accurately perceive the information in the documents from the excessively concatenated images. (4) The performances of different models are highly correlated with their acceptable image numbers and maximum image resolutions. Notably, open-source LVLMs that support high-resolution images (i.e., InternLM-XC2-4KHD and InternVL-Chat-v1.5) exhibit superior performance compared to those with lower resolution limits.
@@
Regarding the human evaluation, we observe 66.0% F1-score from our annotators and a significant performance gap (exceeding 20% in absolute) between the current LVLMs and humans. This gap highlights the challenges of document understanding for LVLMs and the necessity of our benchmark.
@@ 6 Conclusion
Acknowledgements
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). This work is also supported by Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (2022ZD0160201).
@@ A.5 New Question Annotation
A.6 Potential Bias for LVLM-based Quality Checking
As described in Section 3.3, we employ GPT-4o to remove document-agnostic (i.e., can be correctly answer without documents) samples and review potential wrongly-labeled samples. A reasonable speculation raises that our final benchmark can be biased toward GPT-4o’s answers, especially when GPT-4o outperforms others by a large margin. We discuss this potential bias as follows.
We check the effect of GPT-4o’s involvement in the quality control step-by-step. Specifically, we compare the performance of samples remained after each step across GPT-4o and two other competitive models (GPT-4V and Gemini-1.5-Pro). We show their results in the table below.
GPT-4o GPT-4V Gemini-1.5-Pro
No quality control 43.1% 35.2% 23.3%
   + document-relevance detection 41.2% 31.0% 20.5%
   + document-relevance detection + self-reflection / cross-checking 42.7% 31.4% 20.9%
Table 6: Step-wise performance comparison with and without LVLM-based quality checking
The results illustrate that the potential bias in step 1 (document-relevance detection) actually reduce, rather than increase, the performance gap between GPT4o and other models. It is because that we filter out all samples correctly answered by GPT4o without the access to documents. Under this case, the more significant performance drop of GPT-4V and Gemini-1.5-Pro can only be attributed to their limited document understanding and over-reliance on their internal knowledge. Regarding the step 2 and 3 (self-reflection and cross-checking), we provide inconsistent answers between human annotations and GPT4o’s predictions to annotators and ask them to check and revise accordingly. The potential bias of this step does lead to a slight performance bias (1.1% absolute difference at maximum). We believe that such bias is NOT the main cause of GPT4o’s significantly best performance. Without the involvement of GPT-4o in the quality control process, GPT-4o still significantly outperforms GPT-4V by 7.9% (43.1% - 35.2%) and Gemini-1.5-Pro by 19.8% (43.1% - 23.3%). Accordingly, all primary conclusions in our paper still hold.
A.7 GUI Screenshots
@@
A.8 Annotation Cost
This benchmark is annotated by the authors of this paper. Therefore, the data collection does not need compensation. And we count the time cost of our benchmark as below.
Pre-annotation (about 45h): the development of annotation interface (10h), the writing of annotation guideline (5h), training session (10h), preliminary annotation and personalized feedback (20h).
Annotation (about 150h): It takes about 60-90 minutes for the annotation of each document. And all of the 130 documents take about 150 hours.
Post-annotation (about 45h): quality checking (30h), data processing and release preparation (15h).
In summary, our benchmark annotation approximately takes a total of 45+150+45=240 hours (1.36 man months).